OpenStudy (anonymous):

calculate the absolute value of 4-9i

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[\left| x \right|= -x\] and \[x\] So what would \[\left| 4-9i \right|\] be equal to?

OpenStudy (anonymous):

Oh and I meant \[-x \]OR \[x \]not and

OpenStudy (anonymous):

-4+9i

OpenStudy (anonymous):

it's \[-(4+i)\]

OpenStudy (anonymous):

what happens to the 9?

OpenStudy (anonymous):

That was a typo. The 9 remains. But remember, the absolute value equals two values.

OpenStudy (anonymous):

so -4-9i

OpenStudy (anonymous):

that's not how that works

OpenStudy (anonymous):

it's \[\sqrt{4^2+9^2}\]

OpenStudy (anonymous):

it's the square root of the real part squared plus the imaginary part squared.

OpenStudy (anonymous):

sqrt 97?

OpenStudy (anonymous):

Oh yea, he's actually right. It's equal to the absolute value mean distance. But then the absolute property that I told you still holds in other questions.

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

What you said only holds for real x.

OpenStudy (anonymous):

Yes I realised that.

OpenStudy (anonymous):

great

OpenStudy (anonymous):

You would also probably need an equals to something in most cases.